# AI-generated text boundary detection with RoFT

### v2

@@ 1 Introduction
In this work, we experiment with a lesser known dataset called RoFT (Real Or Fake Text), collected by Dugan et al. (2023). Each text in this dataset consists of ten sentences, where the first several sentences are human-written and the rest are machine-generated starting from this prompt, mainly by models from the GPT family (Radford et al., 2019; Brown et al., 2020). We consider several techniques developed for binary ATD, modifying them for this more complex boundary detection setting; e.g., following Tulchinskii et al. (2023a) we adapt intrinsic dimension estimation which is currently considered to be the most robust method for cross-domain and cross-model binary detection. Fig. 1 shows a sample input from RoFT and results of the considered models. Our primary contributions are as follows: (1) we evaluate five approaches to adapt perplexity-based detectors to the task of detecting the boundary between human-written and machine-generated text, discussing the differences in their behavior compared to binary ATD and providing a comprehensive analysis of how perplexity scores react to the machine–human transition in the text; we also compare several backbones for the perplexity evaluation to show which model works the best for this purpose; (2) we present evidence that perplexity scores obtained from small language model, trained on data generated by bigger LLM, provide strong features for boundary detection models, leading to relatively high accuracy of detection and cross-domain robustness related to other detection methods; (3) we introduce and evaluate two ways to adapt the classifiers based on intrinsic dimension estimation for this task, showing how the time series analysis can be used for extracting useful information from Transformer representations; (4) we show how the robustness of boundary detectors (including the fully-tuned RoBERTa baseline) to domain shift depends on the particular properties of the domains; we study the properties of the dataset itself and their effect on the performance of every approach; (5) we enrich the RoFT dataset with GPT-3.5-turbo [^1] https://platform.openai.com/docs/model-index-for-researchers (ChatGPT) generation samples; we share this new dataset with the community, establish baselines, and analyze the behavior of our detectors on it.
@@
Figure 1: Sample input from the ROFT-chatgpt dataset colored according to the perplexities of Phi-1.5 on each sentence; “greener” text corresponds to higher mean perplexity, “blue-er”, to higher standard deviation of the perplexity, red text shows where both mean and std are low. The prompt is above the dotted line, the rest is generated.
@@ 2 Related Work
Artificial text detection (ATD) is a well-studied task with plenty of already considered approaches; see, e.g., the surveys by Yang et al. (2023); Uchendu et al. (2023); Wu et al. (2023a). In our work, we mostly concentrate on adapting two groups of methods to artificial text boundary detection: methods based on perplexity estimation since they are the most widely used and methods based on topological data analysis (TDA) since they have shown promising robustness to domain shift and model shift (see below).Perplexity-based methods are based on the idea that human-written text tends to be more unpredictable and varied compared to AI-generated text which may follow more predictable patterns, i.e., has a lower perplexity score; among them we note GPTZero (Weber-Wulff et al., 2023), Sniffer (Li et al., 2023a), SeqXGPT (Wang et al., 2023), and LLMDet (Wu et al., 2023b). Topological data analysis (TDA) for ATD is inspired by the results of Tulchinskii et al. (2023a) who show that for many artificial text generators the intrinsic dimensions (PHD) of RoBERTa embeddings of the texts created by these generators are typically smaller than those of the texts written by humans; this work has been extended by Kushnareva et al. (2021), and TDA has also proven to be useful for closely related tasks of fake news detection (Tudoreanu, ), authorship attribution (Elyasi & Moghadam, 2019), and detection of synthetic speech (Tulchinskii et al., 2023b).
@@
Other related work. Zeng et al. (2023) used a TriBERT-based approach for artificial text boundary detection in student essays for educational purposes, but with no cross-domain problem setting. Wang et al. (2023) address mixed human and AI generations by sentence-level classification, but they do it with a white-box detector that needs access to the generator model, which severely limits applicability. Recently, a SemEval workshop [^3] https://aclanthology.org/venues/semeval/ introduced a small artificial text boundary detection dataset, namely subtask C of Task 8444https://github.com/mbzuai-nlp/SemEval2024-task8. The subtask asks to predict the word where human text ends and artificial text begins. However, to this moment there is no information on (a) which model was used to generate the samples, (b) what are the particular domains or topics of the generations, and (c) any human baselines. Existing technical reports on the competition (Spiegel & Macko, 2024; Datta et al., 2024; Rad et al., 2024) concentrate on other subtasks of Task 8 such as binary ATD and LLM-authorship attribution, but not the boundary detection subtask.
@@ 3 Approach
Topological Time Series (TTS). Inspired by Tulchinskii et al. (2023a), we explore the potential of topological features based on intrinsic dimensionality (ID); we provide an introduction to topological data analysis (TDA), including the definitions of features, in Appendix A. We hypothesize that geometric variations in token sequences can help identify AI-generated text, so we introduce models that process TDA-based features treating them as time series. For every text, we slide a window of $H=20$ tokens (step size $S=5$) over RoBERTa token embeddings and find the intrinsic dimension (PHD) of the points within the window, as shown in Fig. 4 (Schweinhart, 2020). The time series are then classified with a multi-label SVM with the global alignment kernel (GAK) (Cuturi, 2011), where label correspond to a number of the first generated sentence; this method is called “PHD + TS ML” in the tables.
Topological binary classification. In this approach, we train a binary classifier to distinguish between fake and natural text atop a specific predictor. For the base predictor, we employ intrinsic dimensionality calculated over a sliding window of $20$ tokens. The TLE (tight local) intrinsic dimension estimator was chosen due to its robust performance on small data samples (Amsaleg et al., 2019). For the base classifier, we use gradient boosting trees (Friedman, 2000; 2002). To translate probabilities predicted with a binary classifier into a specific boundary that separates real and fake text, we determine the final label by maximizing $\mathbf{y}\mapsto\operatorname*{arg\,max}_{I\in\mathcal{I}}s_{I}(\mathbf{y},\mathbf{x})$, where the score function $s_{I}$ is defined as $s_{I}(\mathbf{y},\mathbf{x})=-\sum_{j}^{n}I^{j}_{\mathbf{y}}\log\hat{p}(\mathbf{x}^{j})$, and $\hat{p}(\mathbf{x})$ are predictions from the base binary classifier. In this context, we use a binary indicator set $I\in\mathbf{1}^{n}$ consisting of a chunk of $k$ zeros followed by another chunk of $n-k$ ones. This method is called “TLE + TS Binary” in the tables.
@@ 4 Experimental evaluation
Artificial text boundary detection. Table 1 presents the main results of our experiments on the RoFT and RoFT-chatgpt datasets; the majority class prediction is the last ($9$th) class of fully human-written texts. On the original RoFT, we also include the human baseline and the best result reported by Cutler et al. (2021), obtained by a classifier built upon concatenated SRoBERTa embeddings of each sentence. We report the accuracy (Acc) and two metrics that account for the sequential nature of boundary detection: soft accuracy, i.e., the percentage of predictions that differ from the correct label by at most one (SoftAcc1; indeed, Fig. 18 in Appendix F shows that many misclassifications point to an adjacent class, which may be acceptable in real world applications), and mean squared error (MSE).
@@ 5 Discussion and analysis of the results
Detection models. We have found that although RoBERTa-based classifiers demonstrate excellent results for in-domain classification, they lose to perplexity-based methods when tested on texts with new styles and topics not present in the training set. For cross-model generalization, all methods demonstrate insufficient generalization abilities, especially when tested on generations of much larger, but also much smaller models. This suggests that AI content detectors can be fooled by either too good or too bad artificial texts. For perplexity-based classifiers, we have found that the size of the base model is not the main requirement for a good and robust detector. The training dataset for perplexity estimation is more important; e.g., ChatGPT generations were best detected by a much smaller model pretrained on synthetic data generated by the same model family. The above observations suggest that pre-training of the perplexity estimators on a large amount of synthetic data from different generators may become the best choice for AI-generated content detection. We leave it for future research. Topological classifiers demonstrated high robustness over domain transfer, but not very high overall accuracy. Geometric properties of the embeddings for both RoFT and RoFT-chatgpt, however, show the difference between PHD distributions of real and fake RoFT text for different models (Fig. 12) and topics (Figs. 12, 12 in Appendix E). TLE dimension distributions for sentences in human-written and AI-generated parts of the texts are different as well (Figs. 15, 15, 15, Appendix E). This means that topological classifiers may help in robust artificial content detection in some particular domains and models.
Data properties. Next, we summarize the observed properties of the data that influence detection quality and create difficulties for different types of artificial content detectors. (1) First, the length of sentences seems to play an important role, deceiving our classifiers. There is indeed a significant difference between distributions of sentence lengths written by humans and generated by LLMs (see Fig. 3 and Figs. 66 in Appendix E). This is supported by our experiments with the length-based classifier, which sometimes outperforms other methods in terms of accuracy (Tables  1 and 2), but fails to generalize. (2) Label distributions(see Figs. 9 and 9 in Appendix E) vary significantly across models. GPT-2 is especially different from natural texts, and the behavior of our models in cross-model transfer to GPT-2 (Appendix D), suggests that the only model stable under the label distribution shift is perplexity-based regression. Although its accuracy numbers are low, it outperforms the human level in terms of MSE even in out-of-domain evaluation. (3) Special text structure can hurt ATD: e.g., recipes often contain numbered lists, and the first sentence tends to be comparatively long, and as a result the Recipes topic is hard for all classifiers to generalize to. The RoBERTa classifier loses the most here, so we investigated its confusion matrices; Fig. 3 shows that RoBERTa tends to classify all examples in the Recipes topic as fully generated. This kind of text structure can be viewed as an adversarial example for AI-generated texts, and indeed, ChatGPT is known for its tendency to generate structured output with bullet lists. (4) Semantic and grammar inconsistencies. Surprisingly, cross-model generalization of RoBERTa-based and perplexity-based classifiers fails not only for the best OOD generator (davinci), but also for the worst generators: small-size GPT-2 and baseline. Perplexity distributions for texts generated by baseline or GPT-2 are very different from the distributions of texts generated by other models (see Fig. 16). This might explain the poor performance of decision trees and linear classifiers when these generators are excluded from the training set. On the other hand, both generators are among the easiest for human raters, because of the high level of semantic inconsistencies and grammar issues. It seems that the supervised classifier discovers non-regularities in the text as a feature of human-generated texts; indeed, they present in our training set as a part of first-person narratives, especially in the Short Stories dataset. (5) Discourse structure. Finally, analysing the poor generalization to narrative domains, we discover a subtle property of human-generated texts that influences boundary detection, namely the underlying discourse structure of the story that can be measured as peaks of perplexity along the story (Fig. 18, Appendix E). Human-generated texts tend to have 2-3 such peaks, e.g. a peak in the middle of the story corresponding to a plot twist or change of narrative focus. This leads to the difficulties in perplexity-based boundary detection: the perplexity drop in the beginning of the fake text part is confused with the same drop due to the story discourse change. Strong perplexity-based classifiers can handle this issue due to the std feature reflecting token-level perplexity deviation (Fig. 18, Appendix E). Fig. 3b shows that the RoBERTa-based classifier is prone to this discourse-related issue: it often incorrectly predicts the real-fake boundary in the middle of a human-generated story (5-6th sentence). A full analysis of feature distributions across domains and models is given in the Appendix.
@@ Appendix A Intrinsic Dimension Estimation Methods
There is the prompt, used for gpt-3.5-turbo and gpt-4 zero-shot experiment:
@@ Appendix C Cross-domain transfer
Tables 5 and 4 supplement Table 2 from the original text, reporting the results of cross-domain transfer for our methods on the RoFT-chatgpt dataset between four text topics presented in the data. We report in-domain and out-of-domain accuracy: the IN column shows results from domains seen during training, while the OUT column reflects the model’s ability to detect artificial texts in the unseen domain corresponding to this column. For each model, training was done on three out of the four domains, and the resulting model was tested on the fourth, unseen domain; we used $60\%$ of these subsets, mixed together, as the training set, $20\%$ as the validation set, and $20\%$ as the test set for in-domain evaluation.
@@ Appendix E Additional Dataset Analysis
Figure 10: PHD distributions for the real and fake parts of the RoFT dataset by generator models
Figure 11: PHD distributions for the real and fake parts of the RoFT dataset by topics
Figure 12: PHD distributions for the real and fake parts of the RoFT-chatgpt dataset text by topics

### v3

@@ 1 Introduction
In this work, we experiment with a lesser known dataset called RoFT (Real Or Fake Text), collected by Dugan et al. (2023). Each text in this dataset consists of ten sentences, where the first several sentences are human-written and the rest are machine-generated starting from this prompt, mainly by models from the GPT family (Radford et al., 2019; Brown et al., 2020). We consider several techniques developed for binary ATD, modifying them for this more complex boundary detection setting; e.g., following Tulchinskii et al. (2023a) we adapt intrinsic dimension estimation which is currently considered to be the most robust method for cross-domain and cross-model binary detection. Fig. 1 shows a sample input from RoFT and results of the considered models. Our primary contributions are as follows: (1) we evaluate five approaches to adapt perplexity-based detectors to the task of detecting the boundary between human-written and machine-generated text, discussing the differences in their behavior compared to binary ATD and providing a comprehensive analysis of how perplexity scores react to the machine–human transition in the text; we also compare several backbones for the perplexity evaluation to show which model works the best for this purpose; (2) we present evidence that perplexity scores obtained from small language model, trained on data generated by bigger LLM, provide strong features for boundary detection models, leading to relatively high accuracy of detection and cross-domain robustness related to other detection methods; (3) we introduce and evaluate two ways to adapt the classifiers based on intrinsic dimension estimation for this task, showing how the time series analysis can be used for extracting useful information from Transformer representations; (4) we show how the robustness of boundary detectors (including the fully-tuned RoBERTa baseline) to domain shift depends on the particular properties of the domains; we study the properties of the dataset itself and their effect on the performance of every approach; (5) we enrich the RoFT dataset with GPT-3.5-turbo [^1] https://platform.openai.com/docs/model-index-for-researchers (ChatGPT) generation samples; we share this new dataset with the community, establish baselines, and analyze the behavior of our detectors on it. [^2] Our official repository: https://github.com/SilverSolver/ai_boundary_detection.git
@@
Figure 1: Sample input from the ROFT-chatgpt dataset. The rightmost part of the picture shows the true bondary between human-written and machine-generated text. Other three parts are colored according to the features, extracted by various classifiers. Here, gpt3.5 predicts that only two last sentences are generated; TLE + TS binary predicts that generation starts from the second sentence; phi1.5 perplexity predicts that generation starts from the third sentence which is correct answer. See Section  3 for detailed explanation of all the methods and Appendix G for more examples from the dataset.
@@ 2 Related Work
Artificial text detection (ATD) is a well-studied task with plenty of already considered approaches; see, e.g., the surveys by Yang et al. (2023); Uchendu et al. (2023); Wu et al. (2023a). In our work, we mostly concentrate on adapting two groups of methods to artificial text boundary detection: methods based on perplexity estimation since they are the most widely used and methods based on topological data analysis (TDA) since they have shown promising robustness to domain shift and model shift (see below).Perplexity-based methods are based on the idea that human-written text tends to be more unpredictable and varied compared to AI-generated text which may follow more predictable patterns, i.e., has a lower perplexity score; among them we note GPTZero (Weber-Wulff et al., 2023), Sniffer (Li et al., 2023a), SeqXGPT (Wang et al., 2023), and LLMDet (Wu et al., 2023b). Topological data analysis (TDA) for ATD is inspired by the results of Kushnareva et al. (2021) and Tulchinskii et al. (2023a). The latter has shown that for many artificial text generators the intrinsic dimensions (PHD) of RoBERTa embeddings of the texts created by these generators are typically smaller than those of the texts written by humans; also TDA has proven to be useful for closely related tasks of fake news detection (Tudoreanu, ), authorship attribution (Elyasi & Moghadam, 2019), and detection of synthetic speech (Tulchinskii et al., 2023b).
@@
Other related work. Zeng et al. (2023) used a TriBERT-based approach for artificial text boundary detection in student essays for educational purposes, but with no cross-domain problem setting. Wang et al. (2023) address mixed human and AI generations by sentence-level classification, but they do it with a white-box detector that needs access to the generator model, which severely limits applicability. Recently, a SemEval workshop [^4] https://aclanthology.org/venues/semeval/ introduced a small artificial text boundary detection dataset, namely subtask C of Task 8555https://github.com/mbzuai-nlp/SemEval2024-task8. The subtask asks to predict the word where human text ends and artificial text begins. However, to this moment there is no information on (a) which model was used to generate the samples, (b) what are the particular domains or topics of the generations, and (c) any human baselines. Existing technical reports on the competition (Spiegel & Macko, 2024; Datta et al., 2024; Rad et al., 2024) concentrate on other subtasks of Task 8 such as binary ATD and LLM-authorship attribution, but not the boundary detection subtask.
@@ 3 Approach
Topological Time Series (TTS). Inspired by Tulchinskii et al. (2023a), we explore the potential of topological features based on intrinsic dimensionality (ID); we provide an introduction to topological data analysis (TDA), including the definitions of features, in Appendix A. We hypothesize that geometric variations in token sequences can help identify AI-generated text, so we introduce models that process TDA-based features treating them as time series. For every text, we slide a window of $H=100$ tokens (step size $S=1$) over RoBERTa token embeddings and find the intrinsic dimension (PHD) of the points within the window, as shown in Fig. 4 (Schweinhart, 2020). Window size $100$ was chosen as a minimal size where PHD estimator stayed stable. The time series are then classified with a multi-label SVM with the global alignment kernel (GAK) (Cuturi, 2011), where label correspond to a number of the first generated sentence; this method is called “PHD + TS ML” in the tables.
Topological binary classification. In this approach, we train a binary classifier to distinguish between fake and natural text atop a specific predictor. For the base predictor, we employ intrinsic dimensionality calculated over a sliding window of $H=20$ tokens (step size $S=5$) [^6] Stride size 5 was chosen empirically after the preliminary experiments. In these preliminary experiments we found out that for smaller stride sizes the series became too long for the Time Series SVM algorithm, and it wasn’t able to converge.. The TLE (tight local) intrinsic dimension estimator was chosen due to its robust performance on small data samples; window size $20$ was taken directly from Amsaleg et al. (2019), as the authors claim that their algorithm is stable enough “for ‘tight’ localities consisting of as few as 20 sample points”. For the base classifier, we use gradient boosting trees (Friedman, 2000; 2002). To translate probabilities predicted with a binary classifier into a specific boundary that separates real and fake text, we determine the final label by maximizing $\mathbf{y}\mapsto\operatorname*{arg\,max}_{I\in\mathcal{I}}s_{I}(\mathbf{y},\mathbf{x})$, where the score function $s_{I}$ is defined as $s_{I}(\mathbf{y},\mathbf{x})=-\sum_{j}^{n}I^{j}_{\mathbf{y}}\log\hat{p}(\mathbf{x}^{j})$, and $\hat{p}(\mathbf{x})$ are predictions from the base binary classifier. In this context, we use a binary indicator set $I\in\mathbf{1}^{n}$ consisting of a chunk of $k$ zeros followed by another chunk of $n-k$ ones. This method is called “TLE + TS Binary” in the tables.
@@ 4 Experimental evaluation
Artificial text boundary detection. Table 1 presents the main results of our experiments on the RoFT and RoFT-chatgpt datasets; the majority class prediction is the last ($9$th) class of fully human-written texts. On the original RoFT, we also include the human baseline and the best result reported by Cutler et al. (2021), obtained by a classifier built upon concatenated SRoBERTa embeddings of each sentence. We report the accuracy (Acc) and two metrics that account for the sequential nature of boundary detection: soft accuracy, i.e., the percentage of predictions that differ from the correct label by at most one (SoftAcc1; indeed, Fig. 18 in Appendix F shows that many misclassifications point to an adjacent class, which may be acceptable in real world applications), and mean squared error (MSE). More precisely, $SoftAcc1=\frac{1}{n}\sum^{n}_{i=1}{\mathbbm{1}_{\{|y_{i}-\hat{y_{i}}|\leq 1\}}}$ and $MSE=\frac{1}{n}\sum^{n}_{i=1}(y_{i}-\hat{y}_{i})^{2}$, where $y_{i}$ predictions of the algorithm, $\hat{y_{i}}$ true labels, $n$ amount of the examples.
@@ 5 Discussion and analysis of the results
      Refer to caption       Refer to caption
      Refer to caption       Refer to caption
@@
Detection models. We have found that although RoBERTa-based classifiers demonstrate excellent results for in-domain classification, they lose to perplexity-based methods when tested on texts with new styles and topics not present in the training set. For cross-model generalization, all methods demonstrate insufficient generalization abilities, especially when tested on generations of much larger or much smaller models. This suggests that AI content detectors can be fooled by either too good or too bad artificial texts. For perplexity-based classifiers, we have found that the size of the base model is not the main requirement for a good and robust detector. The training dataset for perplexity estimation is more important; e.g., ChatGPT generations were best detected by a much smaller model pretrained on synthetic data generated by the same model family. The above observations suggest that pre-training of the perplexity estimators on a large amount of synthetic data from different generators may become the best choice for AI-generated content detection. We leave it for future research. Topological classifiers demonstrated high robustness over domain transfer, but not very high overall accuracy. Geometric properties of the embeddings for both RoFT and RoFT-chatgpt, however, show the difference between PHD distributions of real and fake RoFT text for different models (Fig. 12) and topics (Figs. 12, 12 in Appendix E). TLE dimension distributions for sentences in human-written and AI-generated parts of the texts are different as well (Figs. 15, 15, 15, Appendix E). This means that topological classifiers may help in robust artificial content detection in some particular domains and models.
Data properties. Next, we summarize the observed properties of the data that influence detection quality and create difficulties for different types of artificial content detectors.
  1. 1.
    First, the length of sentences seems to play an important role, deceiving our classifiers. There is indeed a significant difference between distributions of sentence lengths written by humans and generated by LLMs (see Fig. 3 and Figs. 66 in Appendix E). This is supported by our experiments with the length-based classifier, which sometimes outperforms other methods in terms of accuracy (Tables  1 and 2), but fails to generalize.
  2. 2.
    Label distributions (see Figs. 9 and 9 in Appendix E) vary significantly across models. GPT-2 is especially different from natural texts, and the behavior of our models in cross-model transfer to GPT-2 (Appendix D), suggests that the only model stable under the label distribution shift is perplexity-based regression. Although its accuracy numbers are low, it outperforms the human level in terms of MSE even in out-of-domain evaluation.
  3. 3.
    Special text structure can hurt ATD: e.g., recipes often contain numbered lists, and the first sentence tends to be comparatively long, and as a result the Recipes topic is hard for all classifiers to generalize to. The RoBERTa classifier loses the most here, so we investigated its confusion matrices; Fig. 3 shows that RoBERTa tends to classify all examples in the Recipes topic as fully generated. This kind of text structure can be viewed as an adversarial example for AI-generated texts, and indeed, ChatGPT is known for its tendency to generate structured output with bullet lists.
  4. 4.
    Semantic and grammar inconsistencies. Surprisingly, cross-model generalization of RoBERTa-based and perplexity-based classifiers fails not only for the best OOD generator (davinci), but also for the worst generators: small-size GPT-2 and baseline. Perplexity distributions for texts generated by baseline or GPT-2 are very different from the distributions of texts generated by other models (see Fig. 16). This might explain the poor performance of decision trees and linear classifiers when these generators are excluded from the training set. On the other hand, both generators are among the easiest for human raters, because of the high level of semantic inconsistencies and grammar issues. It seems that the supervised classifier discovers non-regularities in the text as a feature of human-generated texts; indeed, they present in our training set as a part of first-person narratives, especially in the Short Stories dataset.
  5. 5.
    Discourse structure. Finally, analysing the poor generalization to narrative domains, we discover a subtle property of human-generated texts that influences boundary detection, namely the underlying discourse structure of the story that can be measured as peaks of perplexity along the story (Fig. 18, Appendix E). Human-generated texts tend to have 2-3 such peaks, e.g. a peak in the middle of the story corresponding to a plot twist or change of narrative focus. This leads to the difficulties in perplexity-based boundary detection: the perplexity drop in the beginning of the fake text part is confused with the same drop due to the story discourse change. Strong perplexity-based classifiers can handle this issue due to the std feature reflecting token-level perplexity deviation (Fig. 18, Appendix E). Fig. 3b shows that the RoBERTa-based classifier is prone to this discourse-related issue: it often incorrectly predicts the real-fake boundary in the middle of a human-generated story (5-6th sentence).
A full analysis of feature distributions across domains and models is given in the Appendix.
@@ 6 Conclusion
Acknowledgments
The work of Sergey Nikolenko was performed at the Saint Petersburg Leonhard Euler International Mathematical Institute and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-289 dated 06/04/2022).
@@ Appendix A Intrinsic Dimension Estimation Methods
Below we show the prompt used for gpt-3.5-turbo and gpt-4 zero-shot experiments.
@@ Appendix C Cross-domain transfer
Tables 5 and 4 supplement Table 2 from the main text, reporting the results of cross-domain transfer for our methods on the RoFT-chatgpt dataset across four text topics present in the data. We report in-domain and out-of-domain accuracy: the IN column shows results from domains seen during training, while the OUT column reflects the model’s ability to detect artificial texts in the unseen domain corresponding to this column. For each model, training was done on three out of the four domains, and the resulting model was tested on the fourth, unseen domain; we used $60\%$ of these subsets, mixed together, as the training set, $20\%$ as the validation set, and $20\%$ as the test set for in-domain evaluation.
@@ Appendix E Additional Dataset Analysis
Figure 10: PHD distributions for the real and fake parts of the RoFT dataset, by generator models
Figure 11: PHD distributions for the real and fake parts of the RoFT dataset, by topics
Figure 12: PHD distributions for the real and fake parts of the RoFT-chatgpt dataset text, by topics